OpenStudy (anonymous):

A funtion is defined by f(x) = 1/2(10^x + 10^-x) for x in R. Show that: 2f(x)f(y) = f(x+y) + f(x-y)

OpenStudy (anonymous):

you have to write it out and you will see that it works

OpenStudy (anonymous):

that is, write \[f(x+y)+f(x-y)=\frac{1}{2}(10^{x+y}+10^{-(x+y)})+\frac{1}{2}(10^{x-y}+10^{-(x-y)})\] and rearrange the terms

OpenStudy (anonymous):

i don't really mean "rearrange the terms. when you compute \[2f(x)f(y)\] you will get the same thing

OpenStudy (anonymous):

can you show me it? I'm kind of confused..

OpenStudy (anonymous):

let me write it on paper first so i don't make a typo.

OpenStudy (anonymous):

ok really there is nothing to do but multiply \[2f(x)f(y)\] and see that you get the thing i wrote above here goes

OpenStudy (anonymous):

\[2f(x)f(y)=2\times \frac{1}{2}(10^x+10^{-x})\times \frac{1}{2}(10^y+10^{-y})\]

OpenStudy (anonymous):

\[=\frac{1}{2}(10^{x+y}+10^{x-y}+10^{-x+y}+10^{-x-y})\]

OpenStudy (anonymous):

hope it is clear what i did. just like multiplying \[(a+b)(c+d)\] you have to do 4 multiplications and since each one has a base of 10 you are adding the exponents

OpenStudy (anonymous):

in the top line i wrote the right hand side. it is \[f(x+y)+f(x-y)=\frac{1}{2}(10^{x+y}+10^{-(x+y)})+\frac{1}{2}(10^{x-y}+10^{-(x-y)})\]

OpenStudy (anonymous):

now compare exponents and see that you have the same thing exactly

OpenStudy (anonymous):

thank you so much.